OpenStudy (agent_a):

Sulfamic acid (HSO3NH2) is a strong monoprotic acid which can be used to standardize a strong base...

OpenStudy (agent_a):

\[HSO_3NH_2+KOH \rightarrow H_2O+K ^{+}+SO_3NH_2^{+}\] A 0.177-g sample of \[HSO_3NH_2\]required 19.4mL of an aqueous solution of KOH. What is the Molarity of the KOH solution?

OpenStudy (agent_a):

I don't think my solution is right: \[0.177gHSO_3NH_2\times \frac{ 1 mol HSO_3NH_2 }{ 97.1g/mol }=0.001822863028molHSO_3NH_2\] \[\frac{ 0.001822863028molHSO_3NH_2 }{ 19.4mLtimes \frac{ 1L }{ 1000mL } }=0.093962011743MHSO_3NH_2\] and then I realized that I need the molarity of KOH, not \[HSO_3NH_2\] .... So I started with the mass of KOH which is 56.108g KOH, and then I got stuck.

OpenStudy (aaronq):

when you have a monoprotic acid and a base that only abstracts 1 proton (or creates 1 OH), you can use: \(M_{acid}V_{acid}=M_{base}V_{base}\) for your example you'd use \(n_{acid}=M_{base}V_{base}\) instead. or you can work through it the long way (stoichiometrically) to arrive at the same answer.

OpenStudy (agent_a):

But I still need to calculate the Molarity, yes? Because from what I understand, I only know that \[V_2=0.0194L\]and I'd have to solve for the Molarity of the acid and Base as well. I used: \[0.177gHSO_3NH_2 \times \frac{ 1molHSO_3NH_2 }{ 97.1g/molHSO_3NH_2 }=0.001822975673molHSO_3NH_2\] but then I did not proceed because I don't know if I should use 19.4mL of KOH to calculate the molarity of \[HSO_3NH_2\] ....

OpenStudy (aaronq):

you don't need the molarity of the acid, since the moles are equal to the M*V product, you can use what i posted earlier.